# Scattering of a weakly bound dimer from a hard wall in one dimension

### v1

@@ II Integrable cases
Integrable systems are known to preserve chemical composition during scattering processes, even when the incident energy exceeds the dissociation threshold [5] [23]. The model introduced in Sec. I is integrable if the mass ratio $m_{1}/m_{2}$ is 1 or 3, for which the problem can be solved exactly via the Bethe Ansatz [12].
@@
If $m_{1}/m_{2}=3$, the integrability of the model is ensured by a dihedral $D_{6}$ symmetry [14]. Setting $(k_{1},k_{2})=(-\frac{3}{4}K+\frac{\mathrm{i}}{a},-\frac{1}{4}K-\frac{\mathrm{i}}{a})$ and choosing the coefficients $A_{j\pm}$ in Eq. (18) in Ref. [14] to satisfy appropriate boundary conditions, we get the following exact wave function:
@@
Equations (11) and (13) are valid for all $K>0$, and thus the reflection coefficient $R=1$ for all $K$ (even if $K>K_{\text{th}}$) if $m_{1}/m_{2}$ is equal to 11 or 33.
@@ III General Cases
Using the Green’s function method, and assuming that the dissociated part $\psi_{d}(x_{1},x_{2})$ is an outgoing wave at large $X$ if $E>0$, we can express the wave function in Eq. (1) as
$\Psi(x_{1},x_{2})=-\frac{2(1+\beta^{2})}{a\beta}\int_{0}^{\infty}F(x_{1},x_{2}/\beta,z,z/\beta)\psi(z)dz,$ (15)
where $\beta=\sqrt{m_{1}/m_{2}}$, $\psi(z)\equiv\Psi(z,z)$,
$F(y_{1},y_{2},z_{1},z_{2})$ $=F_{0}(\sqrt{(y_{1}-z_{1})^{2}+(y_{2}-z_{2})^{2}})$ (16)
$\quad+F_{0}(\sqrt{(y_{1}+z_{1})^{2}+(y_{2}+z_{2})^{2}})$
$\quad-F_{0}(\sqrt{(y_{1}-z_{1})^{2}+(y_{2}+z_{2})^{2}})$
$\quad-F_{0}(\sqrt{(y_{1}+z_{1})^{2}+(y_{2}-z_{2})^{2}}),$
@@
$F_{0}(r)=\begin{cases}-\frac{1}{2\pi}K_{0}(\sqrt{-2m_{1}E}r),&E<0\\ \frac{1}{2\pi}\ln r,&E=0\\ -\frac{\mathrm{i}}{4}H^{(1)}_{0}(\sqrt{2m_{1}E}r),&E>0\\ \end{cases}.$ (17)
Here $K_{0}(\xi)$ is the decaying Bessel function, and $H_{0}^{(1)}(\xi)$ is the Hankel function of the first kind.
Setting $x_{1}=x_{2}=x$ in Eq. (15), we find that $\psi(x)$ satisfies the following integral equation:
$\psi(x)=-\frac{2(1+\beta^{2})}{a\beta}\int_{0}^{\infty}F(x,x/\beta,z,z/\beta)\psi(z)dz.$ (18)
@@
When the incident kinetic energy of the dimer exceeds its binding energy, such that $K>K_{\text{th}}$, the collision of the dimer with the wall could break up the dimer. We numerically solved Eq. (18) and determined the reflection coefficient $R=|f|^{2}$ for various values of the incident momentum and the mass ratio. We plot $R$ versus $K/K_{\text{th}}$ in Fig. 3 for mass ratios ranging from 1 to 10, and in Fig. 4 for large mass ratios 20, 40, 75.8, and 140. Notably, for mass ratios 1 and 3, the reflection coefficient remains equal to one for all collision energies, indicating full reflection without dissociation, and this is a consequence of integrability discussed in Sec. II. For other mass ratios, $R$ generally first decreases as the collision energy increases, and then increases as the collision energy further increases. Furthermore, systems with larger mass ratios demonstrate a more profound suppression of $R$ for appropriate energies, revealing the effect of mass imbalance on the inelastic scattering process. In Fig. 5 we plot the minimum value of the reflection coefficient, $R_{\text{min}}$, for different mass ratios. The two insets in Fig. 5 show the moderate mass ratio regime $1\leq m_{1}/m_{2}\leq 4.3$ and the regime where $R_{\text{min}}$ is close to zero, respectively. In the moderate mass ratio regime, $R$ remains close to unity, while a significant mass imbalance substantially enhances the probability of dissociation. And we find that $R_{\text{min}}$ can reach zero at a critical mass ratio of about $75.8$ (at $K/K_{\text{th}}\approx 1.24$ as shown in Fig. 4). As the incident momentum goes to infinity, the reflection coefficient eventually approaches unity. This effect can be analytically understood as arising from the asymptotic scaling $R=1-O(1/K^{2})$, which we will derive in Sec. V.
@@ IV Born-Oppenheimer Approximation
is the effective potential experienced by the heavy particle. In Fig. 1 we show the numerical results for the dimer-wall scattering phase shift $\delta$ based on the BO approximation, Eqs. (28), (27), and (29), for the mass ratio $m_{1}/m_{2}=40$ at $K/K_{\text{th}}\leq 1$, and find good agreement with the exact numerical results except when $K/K_{\text{th}}$ is comparable to 1.
@@
where $\Gamma(z)$ is the Gamma function. In Fig. 1 we show the predictions of Eq. (31) for $m_{1}/m_{2}=40$, and find that they are close to our numerical results for the dimer-wall scattering phase shifts at $K\ll K_{\text{th}}$. In Fig. 1 we see that the predictions based on Eqs. (28), (27), and (29) are a better approximation to the exact numerical results for the phase shift than the predictions based on Eq. (31) at $m_{1}/m_{2}=40$, but that the two approximations are both very close to the exact numerical results for the phase shift at $K/K_{\text{th}}\ll 1$.

### v3

@@ I Introduction
Recent experimental progress has enabled the preparation of ultracold heteronuclear molecules with a wide range of mass ratios [2]. For example, mixtures of $^{40}\mathrm{K}$ and $^{6}\mathrm{Li}$ reach a mass ratio of approximately 6.7 [30, 29, 6], while combinations involving heavier fermionic isotopes such as $^{173}\mathrm{Yb}$ with $^{6}\mathrm{Li}$ yield mass ratios up to $\approx 29$ [12, 11]. Mass ratios exceeding these values are not currently achievable with stable atomic species. However, extreme mass ratios could be reached by tuning effective masses in optical lattices [16].
@@ II Integrable cases
Integrable systems are known to preserve chemical composition during scattering processes, even when the incident energy exceeds the dissociation threshold [7, 31]. The model introduced in Sec. I is integrable if the mass ratio $m_{1}/m_{2}$ is 1 or 3. The case $m_{1}/m_{2}=1$ was solved by Bethe ansatz by Gaudin [8]. The ansatz for the mass ratio $m_{1}/m_{2}=3$ is based on a finite superposition of plane waves associated with a dihedral group $D_{6}$ as revealed in [20]. This emergent symmetry is an example of the kinematic symmetries that arise in the “no-tunneling” limit imposed by an impenetrable barrier (the hard wall) as discussed in [13, 14].
@@
If $m_{1}/m_{2}=3$, setting $(k_{1},k_{2})=(-\frac{3}{4}K+\frac{\mathrm{i}}{a},-\frac{1}{4}K-\frac{\mathrm{i}}{a})$ and choosing the coefficients $A_{j\pm}$ in Eq. (18) in Ref. [20] to satisfy appropriate boundary conditions, we get the following exact wave function:
@@
Equations (11) and (13) are valid for all $K>0$, and thus the reflection coefficient $R=1$ for all $K$ (even if $K>K_{\text{th}}$) if $m_{1}/m_{2}$ is equal to 11 or $3$.
@@ III General Cases
We begin by rewriting Eq. (1) in a more convenient form,
$\Big(\frac{\partial^{2}}{\partial y_{1}^{2}}+\frac{\partial^{2}}{\partial y_{2}^{2}}+2m_{1}E\Big)\Psi=-\frac{2(1+\beta^{2})}{a}\delta(y_{1}-\beta y_{2})\Psi,$ (15)
where we have introduced the rescaled coordinates,
$y_{1}=x_{1},\quad y_{2}=\frac{x_{2}}{\beta},\quad\text{with}\quad\beta=\sqrt{\frac{m_{1}}{m_{2}}}.$ (16)
Let $G(y_{1},y_{2},y_{1}^{\prime},y_{2}^{\prime})$ be the Green’s function for the noninteracting problem with hard-wall boundary conditions, satisfying
$\Big(\frac{\partial^{2}}{\partial y_{1}^{2}}+\frac{\partial^{2}}{\partial y_{2}^{2}}+2m_{1}E\Big)G=\delta(y_{1}-y_{1}^{\prime})\delta(y_{2}-y_{2}^{\prime}),$ (17)
with $G=0$ whenever $y_{1}=0$, $y_{2}=0$, $y_{1}^{\prime}=0$ or $y_{2}^{\prime}=0$. This Green’s function can be constructed via the method of images:
$G(y_{1},y_{2},y_{1}^{\prime},y_{2}^{\prime})$ $=G_{0}\Big(\sqrt{(y_{1}-y_{1}^{\prime})^{2}+(y_{2}-y_{2}^{\prime})^{2}}\Big)$ (18)
$\quad+G_{0}\Big(\sqrt{(y_{1}+y_{1}^{\prime})^{2}+(y_{2}+y_{2}^{\prime})^{2}}\Big)$
$\quad-G_{0}\Big(\sqrt{(y_{1}-y_{1}^{\prime})^{2}+(y_{2}+y_{2}^{\prime})^{2}}\Big)$
$\quad-G_{0}\Big(\sqrt{(y_{1}+y_{1}^{\prime})^{2}+(y_{2}-y_{2}^{\prime})^{2}}\Big),$
@@
$G_{0}(r)=\begin{cases}-\frac{1}{2\pi}K_{0}(\sqrt{-2m_{1}E}\,r),&E<0\\ \frac{1}{2\pi}\ln r,&E=0\\ -\frac{\mathrm{i}}{4}H^{(1)}_{0}(\sqrt{2m_{1}E}\,r),&E>0\\ \end{cases}.$ (19)
Here $K_{0}(\xi)$ is the decaying Bessel function, and $H_{0}^{(1)}(\xi)$ is the Hankel function of the first kind. Using the Green’s function method, the wave function in scaled coordinates satisfies the Lippmann–Schwinger equation
$\Psi(\mathbf{y})=-\frac{2(1+\beta^{2})}{a}\iint d^{2}\mathbf{y^{\prime}}G(\mathbf{y},\mathbf{y^{\prime}})\delta(y_{1}^{\prime}-\beta y_{2}^{\prime})\Psi(\mathbf{y^{\prime}}),$ (20)
where $\mathbf{y}=(y_{1},y_{2})$, $\mathbf{y^{\prime}}=(y_{1}^{\prime},y_{2}^{\prime})$, and the integral runs over the first quadrant, $y_{1}^{\prime},y_{2}^{\prime}>0$. Transforming back to the original coordinates $(x_{1},x_{2})$, we obtain
$\Psi(x_{1},x_{2})=-\frac{2(1+\beta^{2})}{a\beta}\int_{0}^{\infty}G(x_{1},x_{2}/\beta,z,z/\beta)\psi(z)dz,$ (21)
where $\psi(z)\equiv\Psi(z,z)$. Setting $x_{1}=x_{2}=x$ in Eq. (21) yields a closed integral equation for $\psi(x)$:
$\psi(x)=-\frac{2(1+\beta^{2})}{a\beta}\int_{0}^{\infty}G(x,x/\beta,z,z/\beta)\psi(z)dz.$ (22)
@@
When the incident kinetic energy of the dimer exceeds its binding energy, such that $K>K_{\text{th}}$, the collision of the dimer with the wall could break up the dimer. We numerically solved Eq. (22) and determined the reflection coefficient $R=|f|^{2}$ for various values of the incident momentum and the mass ratio.
We plot $R$ versus $K/K_{\text{th}}$ in Fig. 3 for mass ratios ranging from 1 to 10, and in Fig. 4 for large mass ratios 20, 40, 75.8, and 140. Notably, for mass ratios 1 and 3, the reflection coefficient remains equal to one for all collision energies, indicating full reflection without dissociation, and this is a consequence of integrability discussed in Sec. II. For other mass ratios, $R$ generally first decreases as the collision energy increases, and then increases as the collision energy further increases. Furthermore, systems with larger mass ratios demonstrate a more profound suppression of $R$ for appropriate energies, revealing the effect of mass imbalance on the inelastic scattering process. In Fig. 5 we plot the minimum value of the reflection coefficient, $R_{\text{min}}$, for different mass ratios. The two insets in Fig. 5 show the moderate mass ratio regime $1\leq m_{1}/m_{2}\leq 4.3$ and the regime where $R_{\text{min}}$ is close to zero, respectively. In the moderate mass ratio regime, $R$ remains close to unity, while a significant mass imbalance substantially enhances the probability of dissociation. And we find that $R_{\text{min}}$ can reach zero at a critical mass ratio of about $75.8$ (at $K/K_{\text{th}}\approx 1.24$ as shown in Fig. 4). As the incident momentum goes to infinity, the reflection coefficient eventually approaches unity. This effect can be analytically understood as arising from the asymptotic scaling $R=1-O(1/K^{2})$, which we will derive in Sec. V.
@@ IV Born-Oppenheimer Approximation
is the effective potential experienced by the heavy particle. In Fig. 1 we show the numerical results for the dimer-wall scattering phase shift $\delta$ based on the BO approximation, Eqs. (31), (32) and (33), for the mass ratio $m_{1}/m_{2}=40$ at $K/K_{\text{th}}\leq 1$, and find good agreement with the exact numerical results except when $K/K_{\text{th}}$ is comparable to 1.
@@
where $\Gamma(z)$ is the Gamma function. In Fig. 1 we show the predictions of Eq. (35) for $m_{1}/m_{2}=40$, and find that they are close to our numerical results for the dimer-wall scattering phase shifts at $K\ll K_{\text{th}}$. In Fig. 1 we see that the predictions based on Eqs. (31), (32) and (33) are a better approximation to the exact numerical results for the phase shift than the predictions based on Eq. (35) at $m_{1}/m_{2}=40$, but that the two approximations are both very close to the exact numerical results for the phase shift at $K/K_{\text{th}}\ll 1$.